OpenStudy (anonymous):

Solve for y: 2/3y + 2 = 6

OpenStudy (anonymous):

\[\frac{2}{3}y +2 = 6\]

OpenStudy (anonymous):

\[\frac{2}{3}y = 6-2 => \frac{2}{3} =4\] \[y = \frac{3}{2} \times 4\] \[y = 6\]

OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)?

OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)?

OpenStudy (anonymous):

What are the domain and range of the relation R: {(-5, -3), (1, 0), (6, 4), (9, 8)}?